# Modeling Annotator Disagreement with Demographic-Aware Experts and Synthetic Perspectives

### v1

@@ 6 Experiment 2: Synthetic Annotations
Experiment 1 showed that DeM-MoE can effectively represent label, especially in settings with dense annotations and rich annotator diversity. However, obtaining such rich annotations across diverse perspectives is often impractical due to cost and logistical constraints. Prior work has suggested that LLMs can approximate group-specific perspectives without training Beck et al. (2024); Hu and Collier (2024). Therefore, we ask could synthetic data help fill this gap for training— could LLMs be used to perform data imputation to better learn demographic perspectives? We hypothesize that (H2) it is possible to achieve moderate alignment with human ratings on subjective tasks using LLM-generated annotations conditioned on demographic personas. Prior to training DeM-MoE with synthetic data, we first assess what capacity LLMs have to model the demographic regularity in our data. This experiment lays the foundation for evaluating how synthetic data complements real annotations in downstream modeling (Experiment 3).
@@ 6.1 Experiment Setup
We evaluate the zero-shot performance [^2] Pilot experiments showed that zero-shot generally performed better than few-shot with our data, we use the former. See Appendix D.2 for details. of four instruction-tuned reasoning LLMs: LLaMA-3.3-70B-Instruct, QwQ-32B, OLMo-2-13B, and Mistral-Nemo-Instruct-2407. As baselines, we include a model that predicts ratings at random, and one that predicts the dataset’s mean rating. Each LLM is told to adopt the perspective of a given demographic persona, provide a short reasoning for the rating, and output the final rating. Full prompt templates are in the Appendix D.1. We evaluate model outputs using Pearson’s $r$ for alignment with human labels, and MAE for accuracy.
@@ 6.2 Results
LLaMA outperforms other models in alignment with human judgments (Table 2). Though the level of alignment varies by task, these results suggest that the model is able to generate predictions that are reasonably aligned in magnitude; an analysis with Pearson’s $r$ (Appendix Table 15) shows the same model trends and confirms that LLM predictions are directionally aligned as well, and reasonably calibrated in magnitude. Overall, LLMs struggle with simulating ratings for conversations (Safety and PCC) tasks that have low human annotator agreement in the first place. MAE is lowest for Safety, but this reflects its narrower 1-3 rating scale rather than higher accuracy; all other datasets uses a 1-5 scale. These findings support our hypothesis (H2): zero-shot demographic prompting helps LLMs to approximate human ratings, and could be useful for data imputation to better learn demographic perspectives. Next, we test whether combining synthetic and real annotations improves performance in data-limited settings.
@@ 7 Experiment 3: Model Training with Real and Synthetic Annotations
The results of Experiment 2 suggest that LLMs can be used to impute moderately-aligned ratings for training. These synthetic data offer several potential benefits: (1) new annotations from underrepresented demographic perspectives, helping reduce bias and improve diversity in training Zhezherau and Yanockin (2024); Chen et al. (2024); Li et al. (2024); and (2) scalable dataset sizes without the cost and time required for additional human labeling Chan et al. (2024); Chung et al. (2022). However, the benefits of synthetic data depend on careful integration. Poorly aligned or noisy synthetic data can introduce bias, harm generalization, and risk misrepresenting minority group perspectives Wyllie et al. (2024); Shumailov et al. (2024); Pereira et al. (2021); Ganev et al. (2022). The optimal method for combining real and synthetic data remains unresolved. Some work prevents model collapse by training on both original and synthetic data Gerstgrasser et al. (2024), while others experiment with pretraining-finetuning schemes or balanced data blending Maini et al. (2024); Zhezherau and Yanockin (2024); Krishna et al. (2021); Doshi et al. (2024). However, these works focus using synthetic data to improve task performance, whereas our work uses the data to improve our ability to model the people labeling for the task.
We present a systematic framework to test configurations of synthetic data generation methods with training strategies to blend real and synthetic data optimally. Our goal is to enhance the performance of DeM-MoE by increasing the diversity of perspectives. We hypothesize that H3: strategic integration of synthetic data into training could improve the task of modeling disagreement, and better representation across various demographics.
@@ 7.1 Experiment Setup
Experiment 3 tests two dimensions of using synthetic data: (1) which synthetic data is generated and (2) how the synthetic data is incorporated during training. We describe each next.
@@ 8 Conclusion
We present DeM-MoE, a demographic-aware mixture of experts model that captures structured variation in annotator disagreement via group-level reasoning patterns. By routing inputs to expert subnetworks based on annotator demographics, DeM-MoE introduces an inductive bias that improves performance on subjective judgments. Across three datasets, it other models in predicting ratings across almost all demographic groups. To address sparse demographic coverage, we evaluate synthetic annotations from zero-shot LLM persona prompting. LLMs are able to reach moderate alignment with human ratings. We also propose blending real and synthetic annotations to improve our task, showing that dataset-specific strategies enhance demographic alignment. These findings promote treating annotator disagreement as a meaningful signal and offer practical methods for scaling perspective-aware learning in NLP.
@@ A.1.1 Data Overview
We give an overview of the five datasets we use in Table. 3
@@
A.1.2 Patient Centered Communication Data
@@
We recruited 594 untrained annotators from the United States on Prolific. Instead of asking the participants directly about the measures, we reworded each to ensure precise definitions (Table 5). We show each snippet to 4 different annotators to capture a variety of opinions. Annotators are shown 15 snippets of different conversations. They are asked to imagine that they are the patient in each snippet, and rate these dimensions of doctor qualities based on what the doctor says in each snippet. After completing the ratings, the annotators are also asked questions about their demographic information, their experience with the medical system, their trust in doctors, and ethnic group-based mistrust. Inter-annotator agreement as measured by Krippendorff’s $\alpha$ ranges from 0.244 to 0.338 depending on the quality dimension (Section 6), with doctor informativeness being the lowest, and doctor warmth being the highest. Cronbach’s $\alpha$=0.958, meaning that although there is a lack of consensus among raters (as perceptions of doctor qualities are highly subjective depending on various factors such as experience with the medical system, or demographic factors), there is high internal consistency—i.e., annotators are likely to consistently give similar scores to similar questions about the same text. We aggregate the nine measurements into three measurements of doctor quality: 1) doctor patient-centered communication (sum of doctor informativeness, supportiveness, and partnership) Street et al. (2007) ; 2) doctor perception of patient communication (sum of patient communication, patient agency, and patient competence); and 3) doctor-patient relational communication (sum of doctor warmth, respect, and empathy) Hovey and Massfeller (2014); Back et al. (2005).
@@
Annotator Past Experience Questions: 1) [hcp freq] During the past 12 months, not counting times you went to an emergency room, how many times did you go to a doctor, nurse, or other health professional to get care for yourself? 333 https://hints.cancer.gov/view-questions/question-detail.aspx?PK_Cycle=1&qid=711
@@ Appendix C Additional Experiment 1 Results
C.2 Experiment 1 within-group expert specialization
@@
Expert usage for each demographic category is shown in Fig. 7 for Politeness, Fig. 8 for Offensiveness, Fig. 9 for Safety, Fig. 10 for Toxicity, and Fig. 11 for PCC.
@@
C.3 Experiment 1 Cross-group expert specialization
@@ Appendix D Additional Experiment 2 Details and Results
D.1 Experiment 2 zero-shot prompts
@@
Patient-Centered Communication Instructions You are an AI assistant skilled at simulating a wide range of demographic perspectives. You are tasked with evaluating snippets of doctor-patient conversations from the perspective of an annotator with the following demographics: {demographics}. Each snippet involves a patient diagnosed with prostate cancer. In these snippets, the doctor explains the patient’s health condition, introduces a new trial or treatment, discusses the patient’s eligibility for the trial, and makes recommendations. Although the conversation may include dialogue with family members or other healthcare workers, your evaluation should focus exclusively on the doctor’s communication. Evaluation Task: Rate the doctor’s performance on the following three qualities: 1. Encourages you to share your opinions 2. Is supportive of you 3. Gives thorough and clear information Rating Scale: 1: Not at all 2: Slightly 3: Moderately 4: Very 5: Extremely
@@
D.2 Zero-shot vs. Few-shot prompting
@@
D.3 Experiment 2: Raw dataset performance
@@
[…]
@@
D.4 Experiment 2: cross-dataset rank consistency

### v3

@@ 6 Experiment 2: Synthetic Annotations
Experiment 1 showed that DeM-MoE can represent labels effectively in settings with dense annotations and diverse annotators. Yet obtaining such data is costly and logistically difficult. Prior work suggests LLMs can approximate group-specific perspectives without training Beck et al. (2024); Hu and Collier (2024). We hypothesize that (H2) LLM-generated annotations conditioned on demographic personas can moderately align with human ratings on subjective tasks. Before training DeM-MoE with synthetic data, we first test LLMs’ ability to capture demographic regularities in our datasets, laying the groundwork for Experiment 3 on combining synthetic and real annotations.
@@ 6.1 Experiment Setup
We evaluate the zero-shot performance [^5] Pilot experiments showed that zero-shot generally performed better than few-shot with our data, we use the former. See Appendix D.3 for details. of four instruction-tuned LLMs with reasoning capabilities: LLaMA-3.3-70B-Instruct, QwQ-32B, OLMo-2-13B, and Mistral-Nemo-Instruct-2407. As baselines, we include a model that predicts ratings at random, and one that predicts the dataset’s mean rating. Each LLM is told to adopt the perspective of a given demographic persona, provide a short reasoning for the rating, and output the final rating. Full prompt templates are in the Appendix D.2. We evaluate using Pearson’s $r$ for alignment with human labels, and MAE for accuracy.
@@ 6.2 Results
LLaMA outperforms other models in alignment with human judgments (Table 2). Though the level of alignment varies by task, these results suggest that the model is able to generate predictions that are reasonably aligned in magnitude; an analysis with Pearson’s $r$ (Appendix Table 26) shows the same model trends and confirms that LLM predictions are directionally aligned as well, and reasonably calibrated in magnitude. Overall, LLMs struggle with simulating ratings for conversations (Safety and PCC) tasks that have low human annotator agreement in the first place. MAE is lowest for Safety, but this reflects its narrower 1-3 rating scale rather than higher accuracy; all other datasets uses a 1-5 scale. To test the construct validity of the synthetic data for rare demographic groups, we conduct a more granular analysis of the performance of different demographic groups relative to their frequency in the data (App. D.1). In general, there isn’t a performance gap between the dominant and minoritized groups (though there are a few exceptions, such as PCC annotators with "Other" race, or Politeness annotators with "less than a high school diploma"). These findings support our hypothesis (H2): zero-shot demographic prompting helps LLMs to approximate human ratings, and could be useful for data imputation to better learn demographic perspectives. Next, we test whether combining synthetic and real annotations improves performance in data-limited settings.
@@ 7 Experiment 3: Model Training with Real and Synthetic Annotations
The results of Experiment 2 suggest that LLMs can be used to impute moderately-aligned ratings for training. These synthetic data offer several potential benefits:1) new annotations from underrepresented demographic perspectives, helping reduce bias and improve diversity in training Zhezherau and Yanockin (2024); Chen et al. (2024); Li et al. (2024); and 2) scalable dataset sizes without the cost and time required for additional human labeling Chan et al. (2024); Chung et al. (2022). However, the benefits of synthetic data depend on careful integration. Poorly aligned or noisy synthetic data can introduce bias, harm generalization, and risk misrepresenting minority group perspectives Wyllie et al. (2024); Shumailov et al. (2024); Pereira et al. (2021); Ganev et al. (2022). The optimal method for combining real and synthetic data remains unresolved. Some work prevents model collapse by training on both original and synthetic data Gerstgrasser et al. (2024), while others experiment with pretraining-finetuning schemes or balanced data blending Maini et al. (2024); Zhezherau and Yanockin (2024); Krishna et al. (2021); Doshi et al. (2024). However, these works focus using synthetic data to improve task performance, whereas our work uses the data to improve our ability to model the people labeling for the task.
We present a systematic framework to test configurations of synthetic data generation methods with training strategies to blend real and synthetic data optimally. Our goal is to enhance the performance of DeM-MoE by increasing the diversity of perspectives. We hypothesize H3: strategic integration of synthetic data into training could improve the task of modeling disagreement, and better representation across various demographics.
@@ 7.1 Experiment Setup
Experiment 3 tests two aspects of using synthetic data: 1)which synthetic data is generated; 2) how the synthetic data is incorporated during training.
@@ 8 Conclusion
We present DeM-MoE, a demographic-aware mixture of experts model that captures structured variation in annotator disagreement through group-level reasoning patterns. By routing inputs to expert subnetworks based on annotator demographics, DeM-MoE introduces an inductive bias that improves performance on subjective judgments. Across three datasets, it outperforms other models in predicting ratings for nearly all demographic groups. To address sparse demographic coverage, we evaluate synthetic annotations from zero-shot LLM persona prompting and find moderate alignment with human ratings. We further blend real and synthetic annotations, showing that dataset-specific strategies enhance demographic alignment. These findings highlight annotator disagreement as a meaningful signal and offer practical methods for scaling perspective-aware learning in NLP.
@@ 9 Limitations
In the context of the PCC data (which we collected via Prolific), there may also be concerns about a biased sampling frame. Although we intentionally increased diversity by recruiting annotators to roughly balance key U.S. demographic groups (e.g., approximately 25% White, 25% Black, 25% Asian, and 25% Other; 50% male, 50% female) and collected detailed demographic information to assess coverage, the sample still reflects the characteristics of Prolific’s online participant pool. Such participants tend to be younger, more educated, and more technologically literate than the general population, and their experiences and expectations of healthcare communication may differ from those of typical patients. Consequently, the distribution of perspectives represented in our PCC annotations may not fully capture variation across less-represented or harder-to-reach populations. This limitation highlights the importance of validating models on datasets drawn from more representative or context-specific populations.
Next, the wording and formatting of annotation tasks could also influence our findings, both for the human annotations in PCC and the LLM-generated annotations in Experiment 2. In PCC, when asking annotators to rate dimensions of doctor communication (e.g., partnership), we frame the question around the concept underlying each construct (e.g., “encourages you to share your opinions”) rather than the construct label itself. This helps standardize annotators’ definitions and understanding based on the literature. However, providing only the construct name and relying on annotators’ folk understanding could yield different yet insightful results, potentially offering a closer simulation of how patients naturally interpret and evaluate physician communication in real encounters. For the LLM annotations, we kept prompts concise across tasks, but some models may benefit from more elaborated instructions. Systematically testing alternative wordings and formats could strengthen confidence in the LLM-generated annotations and improve downstream modeling using these synthetic data (Experiment 3).
@@ A.1.1 Data Overview
We give an overview of the five datasets we use in Table. 3.
A.1.2 Demographic Signals
As a proxy for the strength of demgraphic signals, we report the ridge regression coefficients using the demographic group features for each dataset (Tables 4, 5, 6, 7, 8). Safety shows the weakest demographic signal.
@@ A.1.2 Demographic Signals
Table 4: Ridge regression coefficients for Safety (sorted)
Feature Coefficient
race 0.5595
age 0.1248
gender 0.1125
education 0.1047
Table 5: Ridge regression coefficients for Toxicity (sorted)
Feature Coefficient
age_range 2.0564
education 1.1142
race 0.7610
lgbtq_status 0.6686
gender 0.5123
political_affiliation 0.4758
Table 6: Ridge regression coefficients for Politeness (sorted)
Feature Coefficient
race 1.4273
education 0.8380
age 0.7597
occupation 0.6140
gender 0.2023
Table 7: Ridge regression coefficients for Offensiveness (sorted)
Feature Coefficient
age 1.3508
occupation 1.0809
race 0.9350
education 0.3848
gender 0.2556
Table 8: Ridge regression coefficients for PCC (sorted)
Feature Coefficient
age_group 2.5280
gender 1.5378
edu_level 1.0890
race 0.8457
doc_trust_category 0.6098
occupation 0.5888
hcp_freq 0.3029
ethnic_trust_category 0.2986
A.1.3 Patient Centered Communication Data
@@
We recruited 594 untrained annotators from the United States on Prolific. We aimed to increase the diversity of our own annotation data by sampling annotators on Prolific in a way that balanced key U.S. demographic groups (e.g., targeting approximately 25% White, 25% Black, 25% Asian, and 25% Other, and 50% male, 50% female). Instead of asking the participants directly about the measures, we reworded each to ensure precise definitions (Table 10). We show each snippet to 4 different annotators to capture a variety of opinions. Annotators are shown 15 snippets of different conversations. They are asked to imagine that they are the patient in each snippet, and rate these dimensions of doctor qualities based on what the doctor says in each snippet. After completing the ratings, the annotators are also asked questions about their demographic information, their experience with the medical system, their trust in doctors, and ethnic group-based mistrust. Inter-annotator agreement as measured by Krippendorff’s $\alpha$ ranges from 0.244 to 0.338 depending on the quality dimension (Section 11), with doctor informativeness being the lowest, and doctor warmth being the highest. Cronbach’s $\alpha$=0.958, meaning that although there is a lack of consensus among raters (as perceptions of doctor qualities are highly subjective depending on various factors such as experience with the medical system, or demographic factors), there is high internal consistency—i.e., annotators are likely to consistently give similar scores to similar questions about the same text. We aggregate the nine measurements into three measurements of doctor quality: 1) doctor patient-centered communication (sum of doctor informativeness, supportiveness, and partnership) Street et al. (2007) ; 2) doctor perception of patient communication (sum of patient communication, patient agency, and patient competence); and 3) doctor-patient relational communication (sum of doctor warmth, respect, and empathy) Hovey and Massfeller (2014); Back et al. (2005).
@@
Annotator Past Experience Questions: 1) [hcp freq] During the past 12 months, not counting times you went to an emergency room, how many times did you go to a doctor, nurse, or other health professional to get care for yourself?  [^6] https://hints.cancer.gov/view-questions/question-detail.aspx?PK_Cycle=1&qid=711
@@ Appendix C Additional Experiment 1 Results
C.2 Best zero-shot performance
Comparing these results (Table 20, 21, 22, 23, 24) with the other models in Fig. 1, we see that zero-shot LLaMA consistently performs the worst across datasets because it is not optimized for modeling annotator judgments. Unlike PMF, it cannot capture systematic annotator behaviors or regularities, and unlike MBERT, it does not learn dataset-specific mappings from text features to annotator ratings. However, finetuning with LoRA (LoRA-LLaMA) substantially improves performance, bringing the model closer to the other baselines.
Model Zero-Shot Mean Zero-Shot 95% CI LoRA Mean LoRA 95% CI
Gender 0.847 (0.842–0.853) 0.708 (0.690–0.726)
Race 0.918 (0.904–0.931) 0.565 (0.510–0.619)
Age 0.872 (0.865–0.879) 0.757 (0.740–0.775)
Occupation 0.907 (0.894–0.919) 0.691 (0.661–0.720)
Education 0.915 (0.907–0.924) 0.750 (0.732–0.767)
Table 20: Best zero-shot performance on Offensiveness dataset.
Model Zero-Shot Mean Zero-Shot 95% CI LoRA Mean LoRA 95% CI
Gender 0.785 (0.777–0.794) 0.590 (0.574–0.606)
Race 0.818 (0.801–0.835) 0.628 (0.605–0.651)
Age 0.810 (0.802–0.817) 0.612 (0.602–0.622)
Occupation 0.794 (0.788–0.801) 0.590 (0.580–0.600)
Education 0.773 (0.765–0.781) 0.585 (0.574–0.596)
Table 21: Best zero-shot performance on Politeness dataset.
Model Zero-Shot Mean Zero-Shot 95% CI LoRA Mean LoRA 95% CI
Gender 0.874 (0.869–0.879) 0.627 (0.613–0.641)
Race 0.890 (0.884–0.898) 0.644 (0.628–0.660)
Age 0.881 (0.875–0.886) 0.628 (0.622–0.634)
Education 0.845 (0.837–0.854) 0.639 (0.629–0.649)
Table 22: Best zero-shot performance on Safety dataset.
Model Zero-Shot Mean Zero-Shot 95% CI LoRA Mean LoRA 95% CI
Gender 0.898 (0.891–0.905) 0.692 (0.671–0.713)
Race 0.943 (0.940–0.946) 0.804 (0.788–0.820)
Education 0.910 (0.907–0.913) 0.749 (0.737–0.761)
Age range 0.909 (0.904–0.914) 0.855 (0.826–0.884)
Political affiliation 0.921 (0.920–0.922) 0.745 (0.725–0.765)
LGBTQ status 0.949 (0.947–0.951) 0.802 (0.778–0.826)
Table 23: Best zero-shot performance on Toxicity dataset.
Model Zero-Shot Mean Zero-Shot 95% CI LoRA Mean LoRA 95% CI
hcp freq 0.906 (0.898–0.914) 0.841 (0.827–0.855)
edu level 0.980 (0.967–0.992) 0.919 (0.882–0.956)
age group 0.933 (0.924–0.943) 0.839 (0.827–0.851)
gender 0.968 (0.937–1.000) 0.899 (0.846–0.952)
race 0.975 (0.966–0.984) 0.838 (0.828–0.848)
occupation 0.979 (0.967–0.992) 0.916 (0.892–0.940)
doc trust 0.955 (0.947–0.963) 0.856 (0.842–0.870)
ethnic trust 0.969 (0.961–0.978) 0.881 (0.865–0.897)
Table 24: Best zero-shot performance on PCC dataset.
C.3 Capacity vs. Inductive Bias
Model capacity could potentially confound MoE’s increased performance compared to other models. However, we don’t see this to be the case based on the parameter counts (Table 25). We see that actually LoRA-LLaMA, MBERT, and En + Ea Model have much larger parameter sizes compared to MoE. However, despite having fewer parameters, MoE consistently outperforms other models in most demographic categories on Toxicity, Offensiveness, and PCC. Thus we can confidently say that our proposed architecture, rather than the increase in parameter capacity, provides the beneficial inductive bias.
Model Trainable Parameters (M)
PMF 0.22
LoRA-LLaMA 3.43
MBERT 394.78
En + Ea Model 4.80
Jury Learning 0.49
MoE 2.47
Table 25: Trainable parameter counts for all models compared in the main results.
C.4 Experiment 1 within-group expert specialization
@@
Expert usage for each demographic category is shown in Fig. 7 for Politeness, Fig. 8 for Offensiveness, Fig. 9 for Safety, Fig. 10 for Toxicity, and Fig. 11 for PCC. The figures show subgroup-level averages, where we aggregate expert usage across all instances within each demographic category. This averaging naturally produces distributions where multiple experts appear active, even though routing is discrete per instance.
@@
C.5 Experiment 1 Cross-group expert specialization
@@ Appendix D Additional Experiment 2 Details and Results
D.1 Experiment 2 Group MAE Results
To test the construct validity of the synthetic data for rare demographic groups, we conduct a more granular analysis of the performance of different demographic groups relative to their frequency in the data (Fig 17,18, 19, 20,21). In general, there isn’t a performance gap between the dominant and minoritized groups (though there are a few exceptions, such as PCC annotators with "Other" race, or Politeness annotators with "less than a high school diploma").
Figure 17: MAE by demographic group for the Politeness dataset.
Figure 18: MAE by demographic group for the Offensiveness dataset.
Figure 19: MAE by demographic group for the Toxicity dataset.
Figure 20: MAE by demographic group for the Safety dataset.
Figure 21: MAE by demographic group for the PCC dataset.
D.2 Experiment 2 zero-shot prompts
@@
Patient-Centered Communication Instructions You are an AI assistant skilled at simulating a wide range of demographic perspectives. You are tasked with evaluating snippets of doctor‑patient conversations from the perspective of an annotator with the following demographics: {demographics}. Each snippet involves a patient diagnosed with prostate cancer. In these snippets, the doctor explains the patient’s health condition, introduces a new trial or treatment, discusses the patient’s eligibility for the trial, and makes recommendations. Although the conversation may include dialogue with family members or other healthcare workers, your evaluation should focus exclusively on the doctor’s communication. Evaluation Task: Rate the doctor’s performance on the following three qualities: 1. Encourages you to share your opinions 2. Is supportive of you 3. Gives thorough and clear information Rating Scale: 1: Not at all 2: Slightly 3: Moderately 4: Very 5: Extremely
@@
D.3 Zero-shot vs. Few-shot prompting
@@
D.4 Experiment 2: Raw dataset performance
@@
Model OFF POL Safety PCC TOX Random 0.010 0.004 -0.005 -0.031 0.002 Mean Predictor 0 0 0 0 0 LLaMA-3.3-70B 0.473 0.499 0.201 0.230 0.407 OLMo-2-13B 0.381 0.511 0.0616 0.200 0.395 Mistral-Nemo 0.279 0.424 0.0394 0.169 0.298 QwQ-32B 0.388 0.218 0.120 0.246 0.390
@@
D.5 Experiment 2: cross-dataset rank consistency